Oscar Clemente is the manager of Forbes Division of Pitt, Inc., a manufacturer of biotech products. Forbes Division, which has $4.11 million  in assets, manufactures a special testing device. At the beginning of the current year, Forbes invested $5.05 million in automated equipment for test machine assembly. The division’s expected income statement at the beginning of the year was as follows.

 





Sales revenue$16,080,000
Operating costs


Variable
2,190,000
Fixed (all cash)
7,540,000
Depreciation


New equipment
1,590,000
Other
1,300,000
Division operating profit$3,460,000

 

A sales representative from LSI Machine Company approached Oscar in October. LSI has for $4.56 million a new assembly machine that offers significant improvements over the equipment Oscar bought at the beginning of the year. The new equipment would expand division output by 10 percent while reducing cash fixed costs by 5 percent. It would be depreciated for accounting purposes over a three-year life. Depreciation would be net of the $672,000 salvage value of the new machine. The new equipment meets  Pitt's 12 percent cost of capital criterion. If Oscar purchases the new machine, it must be installed prior to the end of the year. For practical purposes, though, Oscar can ignore depreciation on the new machine because it will not go into operation until the start of the next year.

 

The old machine, which has no salvage value, must be disposed of to make room for the new machine.

 

Pitt has a performance evaluation and bonus plan based on residual income. Pitt uses a cost of capital of 12 percent in computing residual income. Income includes any losses on disposal of equipment. Investment is computed based on the end-of-year balance of assets, net book value. Ignore taxes.

 

Required:

a. What is Forbes Division’s residual income if Oscar does not acquire the new machine?

b. What is Forbes Division’s residual income this year if Oscar acquires the new machine?

c. If Oscar acquires the new machine and operates it according to specifications, what residual income is expected for next year?

 

(Enter your answers in thousands of dollars. Negative amounts should be indicated by a minus sign. Round your answers to the nearest whole dollars)
Answer #1

A) 3,460- 0.12* (4,110+ 5,020- 1,590- 1,300)= 2,707.6 or Answer: 2,708


B) step 1) 5,050- 1,590= 3,460

(3,460 (from division operating profit)- 3,360(step 1)) -0.12* (4,110- 1,300+

4,560)= -884.4 Answer for part B


C) revenue 16,080,000* 10%= 1,608,000+ 16,080,000= 17,688,000

variable  2,190,000* 10%= 219,000+ 2,190,000= 2,409,000

fixed  7,540,000* 5%= 377,000- 7,540,000= 7,163,000

equipment  (4,560,000- 672,000)/ 3 year= 1,296,000

other  1,300,000 same as in question

total 17,688,000- 2,409,000- 7,163,000- 1,296,000- 1,300,000= 5,520,000


5,520- 0.12* (4,110- (2* 1,300)+ 4,560- 1,296)= 4,947.12 Answer for part C.